OpenStudy (anonymous):

Please helpp!? :D Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) x + 4y = –16 3x + 2y = 12

OpenStudy (anonymous):

Part 1: You would say something along the lines of "I would graph both linear lines to find the solution to both of the equations which will be where they intercept (where the two lines meet at each other at one point)."

OpenStudy (anonymous):

Part 2: 1) rearrange x + 4y = -16 into terms of x x + 4y = -16 x = -4y - 16 substitute x = -4y - 16 into the second equation(3x + 2y=12) 3(-4y-16) + 2y=12 -12y -48 + 2y = 12 -10y -48 = 12 -10y = 60 y= 60/-10 y=-6 substitute y=-6 into either equation to find x, lets use the equation x + 4y=-16 x + 4(-6) = -16 x - 24 = -16 x = -16 + 24 x =8 Solution is at (8,-6)

OpenStudy (anonymous):

Part 3: The Solution (8,-6) represents where the two linear equations intercept.

OpenStudy (anonymous):

for Part 1 you may have to add a bit more, I'm not sure what else to add but try to add something like this "The point where they intercept will have the same coordinates for both equations"

OpenStudy (anonymous):

I think I have an idea on what to put for the first part :) Thank you soo freaking much! <3 but i have 2 more if you don't mind?

OpenStudy (anonymous):

k sure, I'll see if I can do them.

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 9y = –16 2x + 6y = –16

OpenStudy (anonymous):

would this work for the first part? First I would change each equation in y=mx+b form. Then I would use the y intercept b and the slope of m to graph both lines.

OpenStudy (anonymous):

Part 1: 1) I would identify the like terms x and y 2) They can be minused.  I'll let u do this part 3) Now that I have the y-value (which you will have found in step 2) 4) Substitute the y-value back into either equation to find the x-value. 5) (your solution)